# Kalman Filter Enhanced GRPO for Reinforcement Learning-Based Language Model Reasoning

### v2

@@ 5.4.2 Model Convergence against Process Noise and Measurement Noise
5.4.3 KL Divergence, Normalized Gradients and Reward Baseline Stabilization Analysis
(a) KL divergence between the RL trained model with the reference model.
@@ 5.4.3 KL Divergence, Normalized Gradients and Reward Baseline Stabilization Analysis
We also examine the KL divergence of action probabilities between the RL finetuned model and its reference model in Fig. 4(a). Interestingly, we notice that the KRPO has slightly higher KL divergence compared to GRPO baseline. It makes sense as we observe from Tab. 1, the Llama3.2 model itself is not strong enough to perform ideally on the reasoning tasks, so deviating from the reference model but not too much would give the model more freedom to explore and fetch higher rewards.
For the gradients as Fig. 4, we applied gradient normalization and clipping to stabilize training by limiting the total gradient norm to a fixed $\ell_{2}$-norm threshold for the whole model. But the visualized gradient is before clipping. We can see the KRPO has a higher gradient than GRPO and gradually goes down, as the estimated variance from the Kalman Filter process can highlight the benefits/drawbacks of actions. The optimization signal is stronger and more informative, making it easier to optimize than its baseline model.
@@
5.4.4 Case Study
Question {"type": "Algebra", "question": "If $3437^{4x}=$ 343, what is the value of $7^{4x-3}$", "expected_answer": "1"}
Model Thinking Process
GRPO First, let’s rewrite $3437^{4x}=$ 343 as $7^{4x}=7^{3}$. This is because $343=7^{3}$. By equating the exponents, we get $34x=$ 3. Solving for $x$, we find that $x=\frac{3}{4}$. Therefore, $3437^{4\cdot\frac{3}{4}}=7^{3}=$ 343 . The answer to this problem is $343$.
KRPO To solve the given equation $3437^{4x}=$ 343, we first express both sides with the same base. Since $343=7^{3}$, we can rewrite the equation as $7^{4x}=7^{3}$. This implies $34x=$ 3, so $x=\frac{3}{4}$. Now that we have the value of $x$, we can find $7^{4x-3}$ by substituting $x=\frac{3}{4}$ into the expression: $17^{4\left(\frac{3}{4}\right)-3}=7^{3-3}=7^{0}=$ 1.
We also provide a case study from OpenMath data as Tab. 2 to show the improved reasoning ability of KRPO model over GRPO baseline. From the case, we observe that GRPO can figure out the prior steps, e.g. “This implies $34x=$ 3, so $x=\frac{3}{4}$”, but it stops here and derives the wrong answer 343. On the contrary, the proposed KRPO can get the correct answer 1. For this question, the KRPO can get 12 corrects out of 12 trials, while GRPO can have 3 corrects out of 12. This is because KRPO can better award and punish correct and incorrect answers via a more accurate advantage estimation.
6 Conclusion
In this work, we improve the model performance over GRPO by introducing KRPO, a Kalman filter-based method for accurate advantage estimation. Our approach replaces the static mean with a Kalman filtered estimate that adapts to latent reward dynamics. KRPO maintains the simplicity of GRPO and requires no modification to the policy network or additional learning parameters. Empirical results show that KRPO improves training convergence and policy performance across environments with dynamic and noisy rewards.
Appendix A Time Efficiency
We also present a comparison of the per question training time between GRPO and the proposed KRPO. The Kalman Filter is a non-parametric method. The training time is measured in seconds. According to the results, GRPO requires 2.271±0.5 seconds to complete training, while KRPO takes 2.288±0.5 seconds. The difference in training time is minimal, indicating that the introduction of the Kalman filter for baseline estimation in KRPO does not introduce much computational overhead compared to GRPO. This suggests that the improved estimation in KRPO is achieved with nearly the same training efficiency. For the testing time, the Kalman filter will not be part of the evaluation process so no time burden will be added by it.
Appendix B More Examples of Case Study
Table 3: More case study of KRPO model question answering.
Dataset Data Sample Thinking Process
Arithmetic {"question": "What is -( -985 - -604 )?", "answer": "381"} The calculation process involves: -( (-985 - (-604) ) ) = -( (-985 + 604 ) ) = -( -381 ) = 381
Arithmetic {"question": "What is 9147 + -1895 + -449?", "answer": "6803"} Calculating the expression: 9147 - 1895 = 7252, then 7252 - 449 = 6803
Arithmetic {"question": "( -74 * -25 ) + -34 = ?", "answer": "1816"} First, -74 * -25 = 1850, then 1850 + -34 = 1816
Arithmetic {"question": "Solve -( 427 - 202 ) + 514", "answer": "289"} Let’s calculate this step by step: -( 427 - 202 ) + 514 = -( 225 ) + 514, -( 225 ) = -225, -225 + 514 = 289
OpenMath {"type": "Prealgebra", "question": "Add $81.76$ to $34.587$ and round your answer to the nearest tenth.} To add 81.76 and 34.587, we first add the numbers: 81.76 + 34.587 = 116.347. Then, we round to the nearest tenth, which gives us 116.3.
OpenMath {"type": "Prealgebra", "question": "What is the greatest common factor of 32 and 48", "expected_answer": "16"} To find the greatest common factor (GCF) of 32 and 48, we need to list the factors of each number. The factors of 32 are: 1, 2, 4, 8, 16, 32. The factors of 48 are: 1, 2, 3, 4, 6, 8, 12, 16, 24, 48. The greatest common factor is the largest number that appears in both lists, which is 16.
OpenMath {"type": "Counting & Probability", "question": "Compute $\dbinom{7}{4}$.", "expected_answer": "35"} To compute $\dbinom{7}{4}$, we use the formula for combinations: $\dbinom{n}{r}=\dfrac{n!}{r!(n-r)!}$. In this case, $n=7$ and $r=4$. Plugging these values into the formula, we get: $\dbinom{7}{4}=\dfrac{7!}{4!(7-4)!}=\dfrac{7!}{4!3!}=\dfrac{7\times 6\times 5}{3\times 2\times 1}=35$.
Appendix C The Analysis of Leading Statement for Language Model Reasoning
From Tab. 1 and Tab. 3 we find that compared with OpenMath-Instruct data, Arithmetic data questions is not harder but model receives much lower results, especially in the hard arithmetic questions. So we hypothesis that the leading statement might be a reason as there are leading statement in OpenMath-Instruct data, but not in Arithmetic data. So we add a leading statement "Please calculate the following expression: " to each arithmetic question.
Interestingly, by enhancing with the leading statement (denoted with KRPO*), on the hard question set, the KRPO performance has increased largely (from 20.881% of KRPO to 40.554% of KRPO*). However, for easy and normal question sets, there are improvements with leading statements but this advantage is not as obvious as on hard question set: on easy set, KRPO 71.764% V.S. KRPO* 72.478%; on the normal set, KRPO 50.589% V.S. 51.403%. It may be caused by the learning saturation of model on easy and normal question sets as the questions have fewer terms and digits that are not that difficult to be handled; but for hard question set with a lot more terms and digits, the model will easily get lost in the understanding of just arithmetic questions, so some leading words can help greatly.

### v3

@@ Appendix G The Analysis of Leading Statement for Language Model Reasoning
Appendix H KL Divergence, Normalized Gradients and Reward Baseline Stabilization Analysis
(a) KL divergence between the RL trained model and the reference model.
@@ Appendix H KL Divergence, Normalized Gradients and Reward Baseline Stabilization Analysis
We also examine the KL divergence of action probabilities between the RL finetuned model and its reference model in Fig. 5(a). Interestingly, we notice that the KRPO has slightly higher KL divergence compared to the GRPO baseline. It makes sense as we observe from Tab. 1, the Llama3.2 model itself is not strong enough to perform ideally on the reasoning tasks, so deviating from the reference model but not too much would give the model more freedom to explore and fetch higher rewards.
For the gradients as Fig. 5, we applied gradient normalization and clipping to stabilize training by limiting the total gradient norm to a fixed $\ell_{2}$-norm threshold for the whole model. But the visualized gradient is before clipping. We can see that the KRPO has a higher gradient than the GRPO and gradually goes down, as the estimated variance from the Kalman Filter process can highlight the benefits/drawbacks of actions. The optimization signal is stronger and more informative, making it easier to optimize than its baseline model.